# LAW 2025: Bridging Language, Agent, and World Models for Reasoning and Planning

### v1

@@ Abstract
We observe that current state-of-the-art web-agents are unable to effectively adapt to new environments without neural network fine-tuning, without which they produce inefficient execution plans due to a lack of awareness of the structure and dynamics of the new environment. To address this limitation, we introduce ATLAS (Actor-Critic Task-completion with Look-ahead Action Simulation), a memory-augmented agent that is able to make plans grounded in a model of the environment by simulating the consequences of those actions in cognitive space. Our agent starts by building a "cognitive map" by performing a lightweight curiosity driven exploration of the environment. The planner proposes candidate actions; the simulator predicts their consequences in cognitive space; a critic analyzes the options to select the best roll-out and update the original plan; and a browser executor performs the chosen action. On the WebArena-Lite Benchmark, we achieve a 63% success rate compared to  53.9% success rate for the previously published state-of-the-art. Unlike previous systems, our modular architecture requires no website-specific LLM fine-tuning. Ablations show sizable drops without the world-model, hierarchical planner, and look-ahead-based replanner confirming their complementary roles within the design of our system.
@@ 1 Introduction
Autonomous agents that can navigate and act on the web have the potential to perform complex tasks like information gathering, transactions, and site configuration on behalf of users (Yao et al., 2022; Sodhi, 2023). However, current web-based agents fall far short of human-level reliability on long-horizon tasks (Koh et al., 2024). The difficulty stems from partial observability, vast action spaces, the need for multi-step planning and memory in a web environment. For example, tasks in WebArena (Zhou et al., 2024) span diverse websites and require reasoning over multi-page navigation and content. An agent might be asked to “Tell me how many fulfilled orders I have over the past three days, and the total amount spent” on an e-commerce site, or “Set the homepage URL on my GitLab profile to https://egg.tart.com.” Solving these tasks demands understanding site structure (e.g. admin dashboards, forms), remembering relevant information (like login states or filters applied), and avoiding irreversible mistakes (like deleting data or purchasing an item unintentionally).
Large language models (LLMs) have markedly improved semantic understanding and generation, suggesting they can enrich web navigation. Yet reliable long-horizon control remains elusive because LLM agents are typically reactive and lack structured memory and explicit planning. Current state-of-the-art web agents such as Plan-and-Act by Erdogan et al. (2025) have significant gaps - the agent’s planner module is not grounded to the structure of the website and requires website-specific model fine-tuning of both planner and executor modules in order to enable new use-cases.
We address these gaps with ATLAS—an inference-time, actor–critic web agent that plans before acting via look-ahead simulation and retrieves structured memories to remain goal-directed over extended interactions. ATLAS has a modular architecture with four components: (1) a Planner that decomposes the task into subgoals; (2) an Actor that proposes diverse next-step candidates; (3) a Critic that forecasts each candidate’s outcome by simulating state transitions and selects the safest, goal-advancing action; and (4) a multi-layer memory that is updated online and queried on demand. Together, these modules perform a simulated look-ahead tree search in conceptual space, enabling adaptive environment-grounded planning, and efficient action selection in realistic, partially observable websites, including those with unexpected environmental hazards.
@@ 2 Related Work
The integration of large language models into autonomous agents has revolutionized web navigation capabilities. Yao et al. (2023) pioneered the ReAct framework, demonstrating how LLMs can effectively combine reasoning and acting in interactive environments. This approach has been extended by Shinn et al. (2024) with Reflexion, which incorporates self-reflection mechanisms for improved decision-making over extended horizons.
@@
Early web navigation systems relied predominantly on rule-based approaches and predefined scripts Liu et al. (2018), which, while interpretable, lacked the adaptability required for dynamic web environments. Recent advances have shifted toward learning-based methods, with several notable developments in autonomous web agents. Zhou et al. (2024) introduced WebArena, a comprehensive benchmark for evaluating web agents across realistic multi-step tasks, establishing a foundation for systematic evaluation in this domain. Building on this, Liu et al. (2024a) introduced WebArena-Lite, a curated subset addressing quality and scalability concerns in the original benchmark.
Recent work has focused on enhancing LLM agents with structured approaches to planning and memory and demonstrated significant progress: Qi et al. (2024) applied reinforcement learning principles to web navigation, achieving notable improvements through policy optimization. Zhang et al. (2025) developed WebPilot, focusing on multimodal understanding of web content, while Erdogan et al. (2025) introduced Plan-and-Act, emphasizing hierarchical task decomposition. Wang et al. (2024b) introduced Agent Workflow Memory (AWM), demonstrating the importance of persistent memory in multi-step web tasks. Yang et al. (2024) developed AgentOccam and demonstrated the effectiveness of simplifying the action space to natural language.
@@
Memory mechanisms have emerged as crucial components for long-horizon task completion. Weston et al. (2014) established foundational work on memory networks, which has been adapted for sequential decision-making contexts. More recently, Zhong et al. (2023) proposed MemoryBank, a comprehensive framework for managing episodic and semantic memory in LLM-based agents. The concept of cognitive maps, originally from cognitive science Tolman (1948), has been adapted for artificial agents. Wayne et al. (2018) demonstrated neural implementations of cognitive mapping in reinforcement learning contexts, while Park et al. (2023) showed how LLMs can maintain and utilize spatial-temporal memory representations for complex behavioral simulation.
@@
Tree search and simulation-based planning have long been central to AI agent design. While classical approaches like Monte Carlo Tree Search Browne et al. (2012) have proven effective in discrete domains, recent work has extended these concepts to natural language environments. Yao et al. (2024) introduced Tree of Thoughts, enabling LLMs to explore multiple reasoning paths through structured search.
World models, which enable agents to simulate future states without environment interaction, have gained renewed attention. Ha & Schmidhuber (2018) established foundational work on learned world models, while more recent efforts by Micheli et al. (2022) have shown how transformer architectures can serve as effective world models for sequential decision-making.
@@
Actor-critic architectures have proven particularly effective for combining policy learning with value estimation. While traditional actor-critic methods focus on reinforcement learning training Sutton & Barto (2018), recent work has adapted these principles to LLM-based agents. The integration of look-ahead planning with actor-critic frameworks has been explored in various contexts, with Silver et al. (2016) demonstrating the power of combining tree search with learned value functions in AlphaGo.
@@
Exploration remains a fundamental challenge in autonomous agent design. Pathak et al. (2017) introduced intrinsic curiosity modules that drive exploration through prediction error, establishing a foundation for self-supervised exploration. Recent work has extended these concepts to language-based environments: Mu et al. (2024) demonstrated curiosity-driven exploration in embodied AI settings.
Our work builds upon these foundations by combining memory-augmented planning with look-ahead simulation in a modular architecture specifically designed for web navigation tasks. Unlike previous approaches that require environment-specific fine-tuning, ATLAS achieves strong performance through inference-time planning and memory retrieval, making it readily adaptable to new domains and underlying models.
@@ 3.2 Architecture Overview
ATLAS comprises four modules operating in an inference-time actor–critic loop with action simulation in conceptual space, shown in Figure 1 (a):
@@
Plans are concise lists of sub-goals with success predicates (e.g., “Reports $\rightarrow$ Sales $\rightarrow$ Set dates $\rightarrow$ Read table”). Planner outputs are included in the context for the actor and critic. Our planner is implemented in the style of (Chae et al., 2025) and extended and described in 3.5.
@@
The utility estimate $V(a)$ is derived via LLM-based assessment that incorporates goal alignment, state viability (recoverability), action coherence, plan consistency, and outcome risk (e.g., destructive or dead-end transitions). Unlike previous efforts that attempt to learn implicit world-model of the environment by fine-tuning a neural network model, such as Chae (2024), we leverage the cognitive map to retrieve action outcomes of each action candidate (Section 3.3). This gives the agent system the ability to look ahead of the current step. We later extend this standard behavior with simulated tree search to further enhance exploratory ability in Section 3.4.
Figure 1: Architecture of ATLAS. (a) Overall flow of ATLAS: The raw observation $o_{t}$ is summarized to lower cognitive load. Then the planner makes a plan $P_{t}$ based on the summarized observation $o^{\prime}_{t}$. The actor proposes NN possible candidate actions for next step. The critic provides judgment of action candidates and finalizes the best action $a_{t}$ to take by considering action outcomes obtained from the cognitive map. (b) Memory construction with curiosity-driven exploration: We build cognitive map by employing exploratory lightweight agents to interact with the environment. (c) Look-ahead Action Simulation (LAS): At each step, ATLAS simulates all candidate actions with the observation from the cognitive map, providing ability to look-ahead. We employ the memory agent to learn from LAS trajectories to make a better plan and update memory if necessary.
@@ 3.3 Memory construction via Curiosity-Driven Exploration
Existing agents may fail because they are 1) not aware of the potential action outcomes, such as placing an order on a shopping website likely to lead to difficulty of canceling and refunding, and, 2) not familiar with environment-specific requirements, such as date format and search format (e.g. AgentOccam-Judge on WebArena-Lite Task 0) (Yang et al., 2024). This is a major gap between LLM agents and human intelligence, since humans can easily predict the outcome of an action with our world knowledge. Encouraging agents to explore the environment and store the findings in memory can effectively avoid actions leading to undesirable outcomes.
@@
Prior work on artificial curiosity Pathak et al. (2017) has demonstrated the ability of neural networks operating in an agentic context to frame curiosity as a self-supervised learning task. Inspired by this insight, we augmented our agent with an artificial curiosity module to initialize the agent. Before evaluation, we perform a curiosity-driven exploration of the web environment to seed the cognitive map and world knowledge of the agent.
@@ 3.5 Look-ahead Action Simulation-backed Dynamic Replanning and Memory Update
As Figure 1 (c) illustrates, the planner integrates a brief exploration digest (what worked/failed, newly exposed affordances, uncovered prerequisites) distilled by the memory writer, then updates $P_{t}$. This can be viewed as a highly simplified implementation of a basic causal learning module that attempts to update our causal model of the world - a highly simplified version of the conceptual flow introduced by Sontakke et al. (2021). This mechanism also prevents catastrophic forgetting of important context, which can happen if the replanner is run with every execution step.
@@ 4 Experimental Setup
WebArena (Zhou et al., 2024) presents a realistic simulation environment comprising a broad array of web navigation tasks such as content retrieval, task execution, and form completion. Tasks vary in complexity, ensuring that the agent’s capabilities are thoroughly tested in realistic scenarios - ranging from purchasing items for eCommerce shopping to updating GitLab code repositories.
The original WebArena consists of 811 tasks, however many of these cannot be performed - humans could only perform 78% of the WebArena. WebArena-Lite is a quality-controlled smaller subset of WebArena consisting of 165 tasks introduced by (Liu et al., 2024b) and has been adapted by prior work in the web agent space such as WebRL (Qi et al., 2024) and Plan-and-Act (Erdogan et al., 2025) as a higher quality and more scalable benchmark for evaluating web agents in the most realistic setting possible - incorporating realistic scenarios such as unexpected environment failures.
@@ 5.1 Comparison with Other Baselines
In this section, we compare our model to other published results on the WebArena-Lite dataset. These are WebPilot (Zhang et al., 2025), Agent Workflow Memory (AWM) (Wang et al., 2024a), WebRL Qi et al. (2024), and Plan-and-Act (Erdogan et al., 2025). Our own work builds on top of AgentOccam (Yang et al., 2024) - we rerun our results on AgentOccam using the state-of-the-art LLM available to us for experimentation at scale, Claude-4 Sonnet.
Table 1: Evaluation Results for ATLAS versus other methods reported on WebArena-Lite. Best performance is in bold
@@ 5.2 Ablation Study
In addition, integrating a high-level planner (Base + HL) in the style of Chae et al. (2025) also improves performance on top of the base agent.
Finally, we construct the final ATLAS agent by integrating both the cognitive map and high-level planner and further extend the system to include replanning via look-ahead search (Base + CM + HL + LA) in order to condition the planner on the cognitive map, we see that the two systems demonstrate complementary performance to produce superior performance in conjunction - achieving state of the art results on WebArena-Lite.
Table 2: Ablation Study Results for Individual Components of ATLAS.
@@ 6 Conclusion
This work presented ATLAS, a web navigation agent that couples explicit, structured memory with hierarchical planning to turn open-ended browsing into a sequence of verifiable, low-entropy decisions. By leveraging contemporary large language models within a modular control loop, the agent maintains situational awareness across pages, decomposes goals into intermediate subgoals, and adapts its strategies as the interface or task constraints evolve. The result is a system that is not only more sample-efficient and time-efficient during exploration, but also more interpretable: intermediate memories, subplans, and decision rationales expose where and why the agent changes course.

### v2

@@ Abstract
Large Language Model (LLM) web agents often struggle with long-horizon web navigation and web task completion in new websites, producing inefficient action sequences unless fine-tuned on environment-specific data. We show that experience-driven memory, combined with look-ahead action simulation, is sufficient for LLM agents to adapt to unseen web environments by remembering past failures and predicting the consequences of future actions. We introduce WebATLAS (Actor-Critic Task-completion with Look-ahead Action Simulation), a memory-augmented LLM web agent that learns a lightweight internal model of the environment from interaction experience and performs hypothetical action rollouts before acting in the real world. WebATLAS builds a persistent cognitive map via curiosity-driven exploration, stores interaction outcomes as experience-based memory, and evaluates candidate actions in cognitive space using a planner–simulator–critic loop. This enables the agent to reuse past experience, avoid previously unsuccessful behaviors, and generate more efficient plans. We evaluate WebATLAS on the WebArena-Lite benchmark for autonomous web navigation and demonstrate a success rate of 63%, outperforming the previous state-of-the-art at 53.9%. Unlike previous systems, our modular architecture requires no website-specific LLM fine-tuning. Ablation studies confirm that experience-driven memory, look-ahead action simulation, and hierarchical replanning play complementary roles in enabling robust, training-free web agents.
@@ 1 Introduction
Autonomous agents that can navigate and act on the web have the potential to perform complex tasks like information gathering, transactions, and site configuration on behalf of users (yao2022webshop; step). However, current web-based agents fall far short of human-level reliability on long-horizon tasks (koh2024tree). The difficulty stems from partial observability, vast action spaces, the need for multi-step planning and memory in a web environment. For example, tasks in WebArena (webarena) span diverse websites and require reasoning over multi-page navigation and content. An agent might be asked to “Tell me how many fulfilled orders I have over the past three days, and the total amount spent” on an e-commerce site, or “Set the homepage URL on my GitLab profile to https://egg.tart.com.” Solving these tasks demands understanding site structure (e.g. admin dashboards, forms), remembering relevant information (like login states or filters applied), and avoiding irreversible mistakes (like deleting data or purchasing an item unintentionally).
Large language models (LLMs) have markedly improved semantic understanding and generation, suggesting they can enrich web navigation. Yet reliable long-horizon control remains elusive because LLM agents are typically reactive and lack structured memory and explicit planning. Current state-of-the-art web agents such as Plan-and-Act by erdogan2025planandact have significant gaps - the agent’s planner module is not grounded to the structure of the website and requires website-specific model fine-tuning of both planner and executor modules in order to enable new use-cases.
We address these gaps with WebATLAS—an inference-time, actor–critic web agent that plans before acting via look-ahead simulation and retrieves structured memories to remain goal-directed over extended interactions. WebATLAS has a modular architecture with four components: (1) a Planner that decomposes the task into subgoals; (2) an Actor that proposes diverse next-step candidates; (3) a Critic that forecasts each candidate’s outcome by simulating state transitions and selects the safest, goal-advancing action; and (4) a multi-layer memory that is updated online and queried on demand. Together, these modules perform a simulated look-ahead tree search in conceptual space, enabling adaptive environment-grounded planning, and efficient action selection in realistic, partially observable websites, including those with unexpected environmental hazards.
@@ 2 Related Work
The integration of large language models into autonomous agents has revolutionized web navigation capabilities. yao2022react pioneered the ReAct framework, demonstrating how LLMs can effectively combine reasoning and acting in interactive environments. This approach has been extended by shinn2024reflexion with Reflexion, which incorporates self-reflection mechanisms for improved decision-making over extended horizons.
@@
Early web navigation systems relied predominantly on rule-based approaches and predefined scripts liu2018reinforcement, which, while interpretable, lacked the adaptability required for dynamic web environments. Recent advances have shifted toward learning-based methods, with several notable developments in autonomous web agents. webarena introduced WebArena, a comprehensive benchmark for evaluating web agents across realistic multi-step tasks, establishing a foundation for systematic evaluation in this domain. Building on this, liu2024vab introduced WebArena-Lite, a curated subset addressing quality and scalability concerns in the original benchmark.
Recent work has focused on enhancing LLM agents with structured approaches to planning and memory and demonstrated significant progress: qi2024webrl applied reinforcement learning principles to web navigation, achieving notable improvements through policy optimization. zhang2025webpilot developed WebPilot, focusing on multimodal understanding of web content, while erdogan2025planandact introduced Plan-and-Act, emphasizing hierarchical task decomposition. wang2024agentworkflowmemory introduced Agent Workflow Memory (AWM), demonstrating the importance of persistent memory in multi-step web tasks. yang2024agentoccam developed AgentOccam and demonstrated the effectiveness of simplifying the action space to natural language.
@@
Memory mechanisms have emerged as crucial components for long-horizon task completion. weston2014memory established foundational work on memory networks, which has been adapted for sequential decision-making contexts. More recently, zhong2024memorybank proposed MemoryBank, a comprehensive framework for managing episodic and semantic memory in LLM-based agents. The concept of cognitive maps, originally from cognitive science tolman1948cognitive, has been adapted for artificial agents. wayne2018unsupervised demonstrated neural implementations of cognitive mapping in reinforcement learning contexts, while park2023generative showed how LLMs can maintain and utilize spatial-temporal memory representations for complex behavioral simulation.
@@
Tree search and simulation-based planning have long been central to AI agent design. While classical approaches like Monte Carlo Tree Search browne2012survey have proven effective in discrete domains, recent work has extended these concepts to natural language environments. yao2024tree introduced Tree of Thoughts, enabling LLMs to explore multiple reasoning paths through structured search.
World models, which enable agents to simulate future states without environment interaction, have gained renewed attention. ha2018world established foundational work on learned world models, while more recent efforts by micheli2022transformers have shown how transformer architectures can serve as effective world models for sequential decision-making.
@@
Actor-critic architectures have proven particularly effective for combining policy learning with value estimation. While traditional actor-critic methods focus on reinforcement learning training sutton2018reinforcement, recent work has adapted these principles to LLM-based agents. The integration of look-ahead planning with actor-critic frameworks has been explored in various contexts, with silver2016mastering demonstrating the power of combining tree search with learned value functions in AlphaGo.
@@
Exploration remains a fundamental challenge in autonomous agent design. pathak2017curiosity introduced intrinsic curiosity modules that drive exploration through prediction error, establishing a foundation for self-supervised exploration. Recent work has extended these concepts to language-based environments: mu2024embodiedgpt demonstrated curiosity-driven exploration in embodied AI settings.
Our work builds upon these foundations by combining memory-augmented planning with look-ahead simulation in a modular architecture specifically designed for web navigation tasks. Unlike previous approaches that require environment-specific fine-tuning, WebATLAS achieves strong performance through inference-time planning and memory retrieval, making it readily adaptable to new domains and underlying models.
@@ 3.2 Architecture Overview
WebATLAS comprises four modules operating in an inference-time actor–critic loop with action simulation in conceptual space, shown in Figure 1 (a):
@@
Plans are concise lists of sub-goals with success predicates (e.g., “Reports $\rightarrow$ Sales $\rightarrow$ Set dates $\rightarrow$ Read table”). Planner outputs are included in the context for the actor and critic. Our planner is implemented in the style of (webshepherd) and extended and described in 3.5.
@@
The utility estimate $V(a)$ is derived via LLM-based assessment that incorporates goal alignment, state viability (recoverability), action coherence, plan consistency, and outcome risk (e.g., destructive or dead-end transitions). Unlike previous efforts that attempt to learn implicit world-model of the environment by fine-tuning a neural network model, such as wma, we leverage the cognitive map to retrieve action outcomes of each action candidate (Section 3.3). This gives the agent system the ability to look ahead of the current step. We later extend this standard behavior with simulated tree search to further enhance exploratory ability in Section 3.4.
Figure 1: Architecture of WebATLAS. (a) Overall flow of WebATLAS: The raw observation $o_{t}$ is summarized to lower cognitive load. Then the planner makes a plan $P_{t}$ based on the summarized observation $o^{\prime}_{t}$. The actor proposes NN possible candidate actions for next step. The critic provides judgment of action candidates and finalizes the best action $a_{t}$ to take by considering action outcomes obtained from the cognitive map. (b) Memory construction with curiosity-driven exploration: We build cognitive map by employing exploratory lightweight agents to interact with the environment. (c) Look-ahead Action Simulation (LAS): At each step, WebATLAS simulates all candidate actions with the observation from the cognitive map, providing ability to look-ahead. We employ the memory agent to learn from LAS trajectories to make a better plan and update memory if necessary.
@@ 3.3 Memory construction via Curiosity-Driven Exploration
Existing agents may fail because they are 1) not aware of the potential action outcomes, such as placing an order on a shopping website likely to lead to difficulty of canceling and refunding, and, 2) not familiar with environment-specific requirements, such as date format and search format (e.g. AgentOccam-Judge on WebArena-Lite Task 0) (yang2024agentoccam). This is a major gap between LLM agents and human intelligence, since humans can easily predict the outcome of an action with our world knowledge. Encouraging agents to explore the environment and store the findings in memory can effectively avoid actions leading to undesirable outcomes.
@@
Prior work on artificial curiosity pathak2017curiosity has demonstrated the ability of neural networks operating in an agentic context to frame curiosity as a self-supervised learning task. Inspired by this insight, we augmented our agent with an artificial curiosity module to initialize the agent. Before evaluation, we perform a curiosity-driven exploration of the web environment to seed the cognitive map and world knowledge of the agent.
@@ 3.5 Look-ahead Action Simulation-backed Dynamic Replanning and Memory Update
As Figure 1 (c) illustrates, the planner integrates a brief exploration digest (what worked/failed, newly exposed affordances, uncovered prerequisites) distilled by the memory writer, then updates $P_{t}$. This can be viewed as a highly simplified implementation of a basic causal learning module that attempts to update our causal model of the world - a highly simplified version of the conceptual flow introduced by causalcuriosity. This mechanism also prevents catastrophic forgetting of important context, which can happen if the replanner is run with every execution step.
@@ 4 Experimental Setup
WebArena (webarena) presents a realistic simulation environment comprising a broad array of web navigation tasks such as content retrieval, task execution, and form completion. Tasks vary in complexity, ensuring that the agent’s capabilities are thoroughly tested in realistic scenarios - ranging from purchasing items for eCommerce shopping to updating GitLab code repositories.
The original WebArena consists of 811 tasks, however many of these cannot be performed - humans could only perform 78% of the WebArena. WebArena-Lite is a quality-controlled smaller subset of WebArena consisting of 165 tasks introduced by (liu2024visualagentbench) and has been adapted by prior work in the web agent space such as WebRL (qi2024webrl) and Plan-and-Act (erdogan2025planandact) as a higher quality and more scalable benchmark for evaluating web agents in the most realistic setting possible - incorporating realistic scenarios such as unexpected environment failures.
@@ 5.1 Comparison with Other Baselines
In this section, we compare our model to other published results on the WebArena-Lite dataset. These are WebPilot (zhang2025webpilot), Agent Workflow Memory (AWM) (wang2024agent), WebRL qi2024webrl, and Plan-and-Act (erdogan2025planandact). Our own work builds on top of AgentOccam (yang2024agentoccam) - we rerun our results on AgentOccam using the state-of-the-art LLM available to us for experimentation at scale, Claude-4 Sonnet.
Table 1: Evaluation Results for WebATLAS versus other methods reported on WebArena-Lite. Best performance is in bold
@@ 5.2 Ablation Study
In addition, integrating a high-level planner (Base + HL) in the style of webshepherd also improves performance on top of the base agent.
Finally, we construct the final WebATLAS agent by integrating both the cognitive map and high-level planner and further extend the system to include replanning via look-ahead search (Base + CM + HL + LA) in order to condition the planner on the cognitive map, we see that the two systems demonstrate complementary performance to produce superior performance in conjunction - achieving state of the art results on WebArena-Lite.
Table 2: Ablation Study Results for Individual Components of WebATLAS.
@@ 6 Conclusion
This work presented WebATLAS, a web navigation agent that couples explicit, structured memory with hierarchical planning to turn open-ended browsing into a sequence of verifiable, low-entropy decisions. By leveraging contemporary large language models within a modular control loop, the agent maintains situational awareness across pages, decomposes goals into intermediate subgoals, and adapts its strategies as the interface or task constraints evolve. The result is a system that is not only more sample-efficient and time-efficient during exploration, but also more interpretable: intermediate memories, subplans, and decision rationales expose where and why the agent changes course.